OpenStudy (anonymous):

A square field has an area of 16km squared. The owner wants to build a fence around the field. How long will the fence be? could you explain how to do it? thanks :)

OpenStudy (saifoo.khan):

16kilometers.

OpenStudy (anonymous):

16km^2= 4km sides. 4*4=16km

OpenStudy (anonymous):

yeah thats what i was gonna do but i wasnt sure. lol thanks

OpenStudy (saifoo.khan):

area of square = L^2 sqrt(16) = L L = 4 Perimter = 4(4) = 16

OpenStudy (anonymous):

thanks everyone :)

OpenStudy (anonymous):

16km * 16km = 256km^2 divide that by four because four sides in a square. Answer is 64km then 64km again divide by four to get 16km